OpenStudy (anonymous):

Use Laplace transform to solve for the IVP y''+2y'+y=0, y(0)=1, y'(0)=1

OpenStudy (anonymous):

s^2*L-s*y(0)-y'(0)+2sL-2*y(0)+L=0 s^2*L-s-1+2s*L-2+L=0 s^2+2s*L+L=s+3 L[s^2+2s+1]=s+3 L=(s+3)/(s+1)^2 s+3 = A/(s+1) + B/(s+1)^2 s+3 = A(s^2+2s+1)+B(s+1) s+3 = As^2+(2A+B)s+A+B 2A+B=1 A+B=3 i get A=-2 and B=5 but the book gets A=1 and A=2 idk where i went wrong